how do I find all points on the y-axis that are 10 units from the point (-6,3)
Answer by stanbon(75887)

how do I find all points on the y-axis that are 10 units from the point (-6,3)
-----------------
Find the equation of all point that are 10 units from(-6,3)
(x+6)^2 + (y-3)^2 = 100
--------------------------
Find the y-intercepts for this equation by letting x=0.
You get:
(0+6)^2 + (y-3)^2 = 100
(y-3)^2 = 64
y-3 = 8 or y-3 = -8
y = 11 or y = -5
====================
The points (0,11) and (0,-5) satisfy the condition.
====================
Cheers,
Stan H.